# StyleDubber: Towards Multi-Scale Style Learning for Movie Dubbing

### v2

@@ Abstract
Given a script, the challenge in Movie Dubbing (Visual Voice Cloning, V2C) is to generate speech that aligns well with the video in both time and emotion, based on the tone of a reference audio track. Existing state-of-the-art V2C models break the phonemes in the script according to the divisions between video frames, which solves the temporal alignment problem but leads to incomplete phoneme pronunciation and poor identity stability. To address this problem, we propose StyleDubber, which switches dubbing learning from the frame level to phoneme level. It contains three main components: (1) A multimodal style adaptor operating at the phoneme level to learn pronunciation style from the reference audio, and generate intermediate representations informed by the facial emotion presented in the video; (2) An utterance-level style learning module, which guides both the mel-spectrogram decoding and the refining processes from the intermediate embeddings to improve the overall style expression; And (3) a phoneme-guided lip aligner to maintain lip sync. Extensive experiments on two of the primary benchmarks, V2C and Grid, demonstrate the favorable performance of the proposed method as compared to the current state-of-the-art. The source code and trained models will be released to the public.
@@
[…]
@@ 1 Introduction
Movie Dubbing Chen et al. (2022), also known as Visual Voice Cloning (V2C), aims to convert a script into speech with the voice characteristics specified by the reference audio, while maintaining lip-sync with a video clip, and reflecting the character’s emotions depicted therein (see Figure 1 (a)). V2C is more challenging than conventional text-to-speech (TTS) Shen et al. (2018a); Ren et al. (2021), and has obvious applications in the film industry and audio AIGC, including broadening the audience for existing video.
@@
According to Zhou et al. (2022); Li et al. (2022), human speech can be perceived as a compound of multi-acoustic factors: (1) unique characteristics, such as timbre, which can be reflected on utterance level (see left panel of Figure 1 (b)); (2) pronunciation habits, such as the rhythm and regional accent, which are usually reflected at the phoneme level (see pink rectangles in Figure 1 (b)). We also note that one’s voice can be affected by emotions. For example, the voice can be significantly different when one gets angry. Based on these observations, we propose to learn phoneme level representations from the speaker’s pronunciation habits reflected in the reference audio, and take both facial expressions and overall timbre characteristics at the utterance level of the reference audio into consideration when generating speech.
@@ 2 Related Work
Text to Speech is a longstanding problem, but recent models represent a dramatic improvement Liu et al. (2024); Tan et al. (2024); Casanova et al. (2022); Wang et al. (2023); Huang et al. (2023a, b). FastSpeech2 Ren et al. (2021), for example, alleviates the one-to-many text-to-speech mapping problem by explicitly modeling variation information. Min et al. (2021), in contrast, improves generalization through episodic meta-learning and generative adversarial networks. Recently Le et al. (2023) proposed a non-autoregressive flow-matching model for mono or cross-lingual zero-shot text-to-speech synthesis. Despite the impressive speech they generate, these methods cannot be applied to the V2C task as they lack the required emotion modelling and lip sync.
Visual Voice Cloning was proposed specifically to address the problem of film dubbing Chen et al. (2022). It requires that a method convert the provided text into speech with the voice and emotion derived from reference audio and a silent video. Building on this, Cong et al. (2023) proposed a hierarchical prosody dubbing model by associating with lip, face, and scene and focus on frame-level prosody learning Hu et al. (2021). To handle multi-speaker scenes, Hassid et al. (2022) match identities by normalizing each speaker to the unit norm and adding in an autoregressive decoder. Recently, Face-TTS Lee et al. (2023) used biometric information extracted directly from the face image as style to improve identity modelling using a score-based diffusion model. Unlike the above methods, StyleDubber address the challenge of insufficient identity information by introducing adaptive utterance-level embedding and detailed pronunciation variations based on the reference audio and video.
Human Pronunciation Modeling aims to learn individual pronunciation variations, which is crucial to generate comprehensible, natural, and acceptable speech Miller (1998). Compared with fixed speaker representations, phoneme-dependent methods Li et al. (2022); Fu et al. (2019) can better control speech and describe more pronunciation features, as phonemes are the basic sound units in a language Lubis et al. (2023). Recently, Zhou et al. (2022) analysed the correlation between local pronunciation content and speaker embeddings at the quasi-phoneme level by reference attention. Here, in contrast, we propose a multimodal style adaptor to capture the fine-grained pronunciation variation, which not only imitates the reference style acoustically, but also conveys emotional expression by reference transformer.
@@ 3.2 Multimodal Phoneme-level Adaptor
Learn acoustic style. We extract reference mel-spectrogram $R_{mel}$ from reference audio $R_{a}$ by Short-time Fourier transform (STFT), and the montreal forced aligner McAuliffe et al. (2017) is used to remove silence phoneme clips. Then, we capture style feature $S_{p}$ via encoder $\mathrm{E_{down}^{spk}}(\cdot)$:
@@
${\mathcal{L}_{spk}=\frac{1}{n}\cdot\sum_{i}^{n}(1-\mathrm{cos\_sim}(\phi(T_{i}),A({S_{p}})_{i})),}$ (5)
where $\phi(\cdot)$ is a function outputting the embedding by the pre-trained G2E2 model Wan et al. (2018), $A(S_{p})$ outputs a style vector via average pooling, and $\mathrm{cos\_sim}(\cdot)$ is the cosine similarity function. $T$ represents the ground truth audio, $n$ is batch size.
Perceive visual emotion. We first use the $S^{3}FD$ model Zhang et al. (2017) to detect facial region from each frame of video, and then an emotion face-alignment network (EmoFAN) Toisoul et al. (2021) is used to extract emotion features $F_{p}\in\mathbb{R}^{{{N_{v}}}\times{D_{m}}}$ from face regions. Similar to style extraction, emotional feature $S_{e}$ is obtained by a downsampling equipped encoder:
@@
Generate intermediate speech representations. We first concatenate the phoneme-level context visual emotion and acoustic style, and then feed it into self-attention blocks $\mathrm{SA}(\cdot)$ to fuse these embeddings:
@@
where $\mathcal{G}_{fus}$ is the fused multimodal context embedding which is with the same length as the textual phonemes. Finally, we combine the textual phoneme embedding and the multimodal context embedding $\mathcal{O}_{pho}=\mathcal{G}_{fus}\oplus T_{H}$, which is viewed as the intermediate dubbing representations.
@@ 3.3 Phoneme-guided Lip Aligner
${{d}_{p}=Total_{Length}\cdot\frac{\mathrm{E_{Softplus}}(C_{lip}^{i})}{\sum_{i}^{N_{p}}\mathrm{E_{Softplus}}(C_{lip})},}$ (11)
@@ 3.5 Training
Dataset V2C-Animation GRID
GT - 100.00 45.58 99.96 0.0 0.0 100.00 22.41 - 0.0 0.0
GT Mel + Vocoder - 96.96 49.33 97.09 3.77 3.80 97.57 21.41 - 4.10 4.15
Fastspeech2 Ren et al. (2021) X 24.87 54.91 42.21 11.20 14.48 47.41 19.05 - 7.67 8.43
StyleSpeech Min et al. (2021) X 54.99 129.67 44.12 11.50 15.10 91.06 24.83 - 5.87 5.98
Zero-shot TTS Zhou et al. (2022) X 48.98 90.59 42.75 9.98 12.51 86.54 19.13 - 5.71 5.99
StyleSpeech* Min et al. (2021) 42.53 130.37 42.53 11.62 14.23 90.04 22.62 - 5.74 5.88
Fastspeech2* Ren et al. (2021) 25.47 54.85 42.39 11.35 14.73 59.58 19.61 - 7.24 7.95
Zero-shot TTS* Zhou et al. (2022) 48.93 87.86 43.97 10.03 12.01 85.93 20.05 - 5.75 6.40
V2C-Net Chen et al. (2022) 40.61 102.69 43.08 14.12 18.49 80.98 47.82 - 6.79 7.23
HPMDubbing Cong et al. (2023) 53.76 187.40 46.61 11.12 11.22 85.11 45.11 - 6.49 6.78
Face-TTS Lee et al. (2023) 52.81 222.31 44.04 13.44 26.94 82.97 44.37 - 7.44 8.16
Ours 82.26 51.82 45.20 9.37 9.46 93.79 18.88 - 5.61 5.69
Table 1: Results under the Dub 1.0 setting Chen et al. (2022), which uses ground-truth audio as reference audio. The method with “*” refers to a variant taking video embedding as an additional input as in Chen et al. (2022). The metric EMO-ACC is not applicable to GRID as it does not have emotional labels.
GT - 100.00 43.59 99.96 0.0 0.0 4.69 $\pm$ 0.12 4.76 $\pm$ 0.09
GT Mel + Vocoder - 96.93 49.11 96.95 3.77 3.80 4.65 $\pm$ 0.07 4.63$\pm$ 0.09
Fastspeech2 Ren et al. (2021) X 24.87 54.91 42.21 11.20 14.48 2.13$\pm$ 0.09 3.75$\pm$ 0.12
StyleSpeech Min et al. (2021) X 42.09 100.01 41.55 11.56 15.10 3.35 $\pm$ 0.07 3.24$\pm$ 0.08
Zero-shot TTS Zhou et al. (2022) X 38.97 81.51 39.11 10.68 13.52 3.58 $\pm$0.11 3.72$\pm$ 0.15
StyleSpeech* Min et al. (2021) 42.52 103.07 42.57 11.58 15.23 3.46 $\pm$ 0.16 3.83$\pm$ 0.15
Fastspeech2* Ren et al. (2021) 25.47 54.85 42.39 11.35 14.73 2.46 $\pm$0.06 3.77 $\pm$0.08
Zero-shot TTS* Zhou et al. (2022) 39.71 82.17 39.30 10.76 13.66 3.68$\pm$ 0.14 3.69 $\pm$0.09
V2C-Net Chen et al. (2022) 30.64 92.90 41.01 14.58 18.73 3.04 $\pm$0.15 2.78 $\pm$0.06
HPMDubbing Cong et al. (2023) 34.10 189.28 43.97 11.88 11.98 3.19 $\pm$0.10 3.06 $\pm$0.22
Face-TTS Lee et al. (2023) 34.14 200.18 43.56 13.78 28.03 3.13$\pm$ 0.12 3.09 $\pm$0.06
Ours 82.24 49.83 45.51 9.44 9.53 4.21 $\pm$0.08 4.05 $\pm$0.09
@@ 4.1 Implementation Details
The video frames are sampled at 25 FPS and all audios are resampled to 22.05kHz. The ground-truth of phoneme duration is extracted by montreal forced aligner McAuliffe et al. (2017). The window length, frame size, and hop length in STFT are 1024, 1024, and 256, respectively. The lip region is resized to 96 $\times$ 96 and pretrained on Resnet18. We use 8 heads for multi-head attention in PLA and the hidden size is 512. Our duration predictor consists of 2-layer 1D convolutional with kernel size 1. The weights in Eq. 16 are set to $\lambda_{1}$ = 20.0, $\lambda_{2}$ = 1.0, $\lambda_{3}$ = 1.0, $\lambda_{4}$ = 2.0, $\lambda_{5}$ = 5.0. For downsampling encoder in $\mathrm{E_{down}^{spk}}$, we use 4 convolutions containing [128, 256, 512, 512] filters with shape 3 $\times$ 1 respectively, each followed by an average pooling layer with kernel size 2. In $\mathrm{E_{down}^{emo}}$, 2 convolutions are used to download to quasi phoneme-level, containing [128, 256] filters with shape 3 $\times$ 1. In $R_{A\rightarrow L}$ and $Z_{V\rightarrow L}$, the dimensionality of all reference attention hidden is set to 128 implemented by a 1D temporal convolutional layer. A well-trained HiFi-GAN is used as vocoder to convert the 80-channel mel-spectrograms to time-domain waveforms. For training, we use Adam Kingma and Ba (2015) with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, $\epsilon$=$10^{-9}$ to optimize our model. The learning rate is set to 0.00625. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU.
@@ 4.2 Dataset
V2C-Animation dataset Chen et al. (2022) is currently the only publicly available movie dubbing dataset for multi-speaker. Specifically, it contains 153 diverse characters extracted from 26 Disney cartoon movies, specified with speaker identity and emotion annotations. The whole dataset has 10,217 video clips, and the audio samples are sampled at 22,050Hz with 16 bits. In practice, Chen et al. (2022) removes video clips less than 1s.
GRID dataset Cooke et al. (2006) is a basic benchmark for multi-speaker dubbing. The whole dataset has 33 speakers, each with 1000 short English samples. All participants were recorded in a noise-free studio with a unified screen background. The train set consists of 32,670 samples, 900 sentences from each speaker. In the test set, there are 100 samples of each speaker.

### v3

@@ Abstract
Given a script, the challenge in Movie Dubbing (Visual Voice Cloning, V2C) is to generate speech that aligns well with the video in both time and emotion, based on the tone of a reference audio track. Existing state-of-the-art V2C models break the phonemes in the script according to the divisions between video frames, which solves the temporal alignment problem but leads to incomplete phoneme pronunciation and poor identity stability. To address this problem, we propose StyleDubber, which switches dubbing learning from the frame level to phoneme level. It contains three main components: (1) A multimodal style adaptor operating at the phoneme level to learn pronunciation style from the reference audio, and generate intermediate representations informed by the facial emotion presented in the video; (2) An utterance-level style learning module, which guides both the mel-spectrogram decoding and the refining processes from the intermediate embeddings to improve the overall style expression; And (3) a phoneme-guided lip aligner to maintain lip sync. Extensive experiments on two of the primary benchmarks, V2C and Grid, demonstrate the favorable performance of the proposed method as compared to the current state-of-the-art. The code will be made available at https://github.com/GalaxyCong/StyleDubber.
@@
Gaoxiang Cong1   Yuankai Qi2111Corresponding authors   Liang Li1111Corresponding authors   Amin Beheshti2   Zhedong Zhang3    Anton van den Hengel4    Ming-Hsuan Yang5   Chenggang Yan3   Qingming Huang1 1Institute of Computing Technology, CAS   2Macquarie University 3Hangzhou Dianzi University   4University of Adelaide   5University of California conggaoxiang@foxmail.com, yuankai.qi@mq.edu.au, liang.li@ict.ac.cn
@@ 1 Introduction
Movie Dubbing Chen et al. (2022a), also known as Visual Voice Cloning (V2C), aims to convert a script into speech with the voice characteristics specified by the reference audio, while maintaining lip-sync with a video clip, and reflecting the character’s emotions depicted therein (see Figure 1 (a)). V2C is more challenging than conventional text-to-speech (TTS) Shen et al. (2018a); Ren et al. (2021), and has obvious applications in the film industry and audio AIGC, including broadening the audience for existing video.
@@
According to Zhou et al. (2022); Li et al. (2022b), human speech can be perceived as a compound of multi-acoustic factors: (1) unique characteristics, such as timbre, which can be reflected on utterance level (see left panel of Figure 1 (b)); (2) pronunciation habits, such as the rhythm and regional accent, which are usually reflected at the phoneme level (see pink rectangles in Figure 1 (b)). We also note that one’s voice can be affected by emotions. For example, the voice can be significantly different when one gets angry. Based on these observations, we propose to learn phoneme level representations from the speaker’s pronunciation habits reflected in the reference audio, and take both facial expressions and overall timbre characteristics at the utterance level of the reference audio into consideration when generating speech.
@@ 2 Related Work
Text to Speech is a longstanding problem, but recent models represent a dramatic improvement Liu et al. (2024); Tan et al. (2024); Casanova et al. (2022); Wang et al. (2023a); Huang et al. (2023a, b); Ju et al. (2024). FastSpeech2 Ren et al. (2021), for example, alleviates the one-to-many text-to-speech mapping problem by explicitly modeling variation information. Min et al. (2021), in contrast, improves generalization through episodic meta-learning and generative adversarial networks. Recently Le et al. (2023) proposed a non-autoregressive flow-matching model for mono or cross-lingual zero-shot text-to-speech synthesis. Despite the impressive speech they generate, these methods cannot be applied to the V2C task as they lack the required emotion modelling and lip sync.
Visual Voice Cloning is proposed to address the problem of film dubbing Chen et al. (2022a) and has attracted a lot of attention in cross-modality alignment field Tu et al. (2022, 2023, 2024); Li et al. (2022a); Liu et al. (2023); Wang et al. (2023b); Xiao et al. (2023, 2022). Then, Cong et al. (2023) proposed a hierarchical prosody dubbing model by associating with lip, face, and scene and focus on frame-level prosody learning Hu et al. (2021). To handle multi-speaker scenes, Hassid et al. (2022) matches identities by normalizing each speaker to the unit norm, and Lu et al. (2022) adopts a lookup table to match the d-vector. Recently, Face-TTS Lee et al. (2023) used biometric information extracted directly from the face image as style to improve identity modelling using a score-based diffusion model. Unlike the above methods, StyleDubber address the challenge of insufficient identity information by introducing adaptive utterance-level embedding and detailed pronunciation variations based on the reference audio and video.
Human Pronunciation Modeling aims to learn individual pronunciation variations, which is crucial to generate comprehensible, natural, and acceptable speech Miller (1998). Compared with fixed speaker representations, phoneme-dependent methods Li et al. (2022b); Fu et al. (2019) can better control speech and describe more pronunciation features, as phonemes are the basic sound units in a language Lubis et al. (2023). Recently, Zhou et al. (2022) analysed the correlation between local pronunciation content and speaker embeddings at the quasi-phoneme level by reference attention. Here, in contrast, we propose a multimodal style adaptor to capture the fine-grained pronunciation variation, which not only imitates the reference style acoustically, but also conveys emotional expression by reference transformer.
@@ 3.2 Multimodal Phoneme-level Adaptor
Learn acoustic style. We extract reference mel-spectrogram $R_{mel}$ from reference audio $R_{a}$ by Short-time Fourier transform (STFT), and the montreal forced aligner McAuliffe et al. (2017) is used to clip phoneme. Then, we capture the style feature $S_{p}$ via a mel-style encoder $\mathrm{E_{down}^{spk}}(\cdot)$:
@@
${\mathcal{L}_{spk}=\frac{1}{n}\cdot\sum_{j}^{n}(1-\mathrm{cos\_sim}(\phi(T_{j}),A({S_{p}})_{j})),}$ (5)
where $\phi(\cdot)$ is a function outputting the embedding by the pre-trained GE2E model Wan et al. (2018), $A(S_{p})$ outputs a style vector via average pooling, and $\mathrm{cos\_sim}(\cdot)$ is the cosine similarity function. $T$ represents the ground truth audio, $n$ is batch size.
Perceive visual emotion. We first use the $S^{3}FD$ model Zhang et al. (2017) to detect facial region from each frame of video, and then an emotion face-alignment network (EmoFAN) Toisoul et al. (2021) is used to extract emotion features $F_{p}\in\mathbb{R}^{{{N_{v}}}\times{D_{m}}}$ from face regions. The $N_{v}$ represents the number of video frames. Similar to style extraction, emotional feature $S_{e}$ is obtained by a downsampling equipped encoder:
@@
Generate intermediate speech representations. We first concatenate the phoneme-level context visual emotion and acoustic style in channel dimension, and then feed it into self-attention blocks $\mathrm{SA}(\cdot)$ to fuse these embeddings:
@@
where $\mathcal{G}_{fus}\in\mathbb{R}^{{{N_{p}}}\times{D_{m}}}$ is the fused multimodal context embedding which is with the same length as the textual phonemes. Finally, we combine textual phoneme embedding and multimodal context embedding $\mathcal{O}_{pho}=\mathcal{G}_{fus}\oplus T_{H}$, which is viewed as the intermediate dubbing representations.
@@ 3.3 Phoneme-guided Lip Aligner
Dataset V2C-Animation GRID
GT - 100.00 22.55 99.96 0.0 0.0 100.00 22.41 - 0.0 0.0
GT Mel + Vocoder - 96.96 24.58 97.09 3.77 3.80 97.57 21.41 - 4.10 4.15
Fastspeech2 Ren et al. (2021) X 24.87 34.48 42.21 11.20 14.48 47.41 19.05 - 7.67 8.43
StyleSpeech Min et al. (2021) X 54.99 106.73 44.12 11.50 15.10 91.06 24.83 - 5.87 5.98
Zero-shot TTS Zhou et al. (2022) X 48.98 68.81 42.75 9.98 12.51 86.54 19.13 - 5.71 5.99
StyleSpeech* Min et al. (2021) 42.53 108.00 42.53 11.62 14.23 90.04 22.62 - 5.74 5.88
Fastspeech2* Ren et al. (2021) 25.47 33.53 42.39 11.35 14.73 59.58 19.61 - 7.24 7.95
Zero-shot TTS* Zhou et al. (2022) 48.93 68.05 43.97 10.03 12.01 85.93 20.05 - 5.75 6.40
V2C-Net Chen et al. (2022a) 40.61 73.08 43.08 14.12 18.49 80.98 47.82 - 6.79 7.23
HPMDubbing Cong et al. (2023) 53.76 164.16 46.61 11.12 11.22 85.11 45.11 - 6.49 6.78
Face-TTS Lee et al. (2023) 52.81 201.13 44.04 13.44 26.94 82.97 44.37 - 7.44 8.16
Ours 82.26 31.49 45.62 9.37 9.46 93.79 18.88 - 5.61 5.69
Table 1: Results under the Dub 1.0 setting Chen et al. (2022a), which uses ground-truth audio as reference audio. The method with “*” refers to a variant taking video embedding as an additional input as in Chen et al. (2022a). The metric EMO-ACC is not applicable to GRID as it does not have emotional labels.
GT - 100.00 22.76 99.96 0.0 0.0 4.69 $\pm$ 0.12 4.76 $\pm$ 0.09
GT Mel + Vocoder - 96.93 24.83 96.95 3.77 3.80 4.65 $\pm$ 0.07 4.63$\pm$ 0.09
Fastspeech2 Ren et al. (2021) X 24.17 35.08 42.21 11.20 14.48 2.13$\pm$ 0.09 3.75$\pm$ 0.12
StyleSpeech Min et al. (2021) X 75.66 76.58 41.55 11.56 15.10 3.35 $\pm$ 0.07 3.24$\pm$ 0.08
Zero-shot TTS Zhou et al. (2022) X 47.79 58.82 39.11 10.68 13.52 3.58 $\pm$0.11 3.72$\pm$ 0.15
StyleSpeech* Min et al. (2021) 75.67 82.48 42.57 11.58 15.23 3.46 $\pm$ 0.16 3.83$\pm$ 0.15
Fastspeech2* Ren et al. (2021) 25.47 34.08 42.39 11.35 14.73 2.46 $\pm$0.06 3.77 $\pm$0.08
Zero-shot TTS* Zhou et al. (2022) 47.55 58.81 39.30 10.76 13.66 3.68$\pm$ 0.14 3.69 $\pm$0.09
V2C-Net Chen et al. (2022a) 34.07 61.61 41.01 14.58 18.73 3.04 $\pm$0.15 2.78 $\pm$0.06
HPMDubbing Cong et al. (2023) 31.42 171.03 43.97 11.88 11.98 3.19 $\pm$0.10 3.06 $\pm$0.22
Face-TTS Lee et al. (2023) 51.98 200.18 43.56 13.78 28.03 3.13$\pm$ 0.12 3.09 $\pm$0.06
Ours 81.27 31.70 41.35 10.59 10.68 3.92 $\pm$0.11 3.86 $\pm$0.09
@@
${{d}_{p}=Total_{Length}\cdot\frac{\mathrm{E_{Softplus}}(C_{lip}^{k})}{\sum_{k=0}^{N_{p}-1}\mathrm{E_{Softplus}}(C_{lip})},}$ (11)
@@ 4.1 Implementation Details
The video frames are sampled at 25 FPS and all audios are resampled to 22.05kHz. The ground-truth of phoneme duration is extracted by montreal forced aligner McAuliffe et al. (2017). The window length, frame size, and hop length in STFT are 1024, 1024, and 256, respectively. The lip region is resized to 96 $\times$ 96 and pretrained on ResNet-18, following Martinez et al. (2020); Ma et al. (2021). We use 8 heads for multi-head attention in PLA and the hidden size is 512. The duration predictor consists of 2-layer 1D convolution with kernel size 1. The weights in Eq. 16 are set to $\lambda_{1}$ = 25.0, $\lambda_{2}$ = 0.1, $\lambda_{3}$ = 5.0, $\lambda_{4}$ = 2.0, $\lambda_{5}$ = 5.0. For downsampling encoder in $\mathrm{E_{down}^{spk}}$, we use 4 convolutions containing [128, 256, 512, 512] filters with shape 3 $\times$ 1 respectively, each followed by an average pooling layer with kernel size 2. In $\mathrm{E_{down}^{emo}}$, 2 convolutions are used to download to quasi phoneme-level, containing [128, 256] filters with shape 3 $\times$ 1. In $R_{A\rightarrow L}$ and $Z_{V\rightarrow L}$, the dimensionality of all reference attention hidden size is set to 128 implemented by a 1D temporal convolutional layer. We set the batch size to 32 and 64 on V2C-Animation and Grid dataset, respectively. For training, we use Adam Kingma and Ba (2015) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, $\epsilon$=$10^{-9}$ to optimize our model. The learning rate is set to 0.00625. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU.
@@ 4.2 Dataset
V2C-Animation dataset Chen et al. (2022a) is currently the only publicly available movie dubbing dataset for multi-speaker. Specifically, it contains 153 diverse characters extracted from 26 Disney cartoon movies, specified with speaker identity and emotion annotations. The whole dataset has 10,217 video clips, and the audio samples are sampled at 22,050Hz with 16 bits. In practice, Chen et al. (2022a) removes video clips less than 1s. In this work, all experiments are conducted on the V2C denoise version. We will publish this version.
GRID dataset Cooke et al. (2006) is a basic benchmark for multi-speaker dubbing. The whole dataset has 33 speakers, each with 1000 short English samples. All participants are recorded in a noise-free studio with a unified screen background. The train set consists of 32,670 samples, 900 sentences from each speaker. In the test set, there are 100 samples of each speaker.